Calculate the critical angle of a material having a refractive index of 1.52 when light passes from the material to air
1
Expert's answer
2019-08-26T10:50:48-0400

Relationship between refractive index and critical angle of a material:

"\mu=\frac{1}{sin \space {\theta}_c}"

where "\theta_c" represent critical angle.

This equation reduces to:

"\theta_c=sin^{-1}\frac{1}{\mu}"

"\theta_c=sin^{-1} \frac{1}{1.52}"

"\theta_c=41.139\degree"